# A deep graph model for the signed interaction prediction in biological network

### v1

@@ Abstract
In pharmaceutical research, the strategy of drug repurposing accelerates the development of new therapies while reducing R&D costs. Network pharmacology lays the theoretical groundwork for identifying new drug indications, and deep graph models have become essential for their precision in mapping complex biological networks. Our study introduces an advanced graph model that utilizes graph convolutional networks and tensor decomposition to effectively predict signed chemical-gene interactions. This model demonstrates superior predictive performance, especially in handling the polar relations in biological networks. Our research opens new avenues for drug discovery and repurposing, especially in understanding the mechanism of actions of drugs.
@@ 1 Introduction
In the new era of pharmaceutical research and development, the strategy of drug repurposing plays a particularly critical role in the exploration of innovative therapies. This approach, through the rediscovery of existing drugs for new pathways or diseases, not only accelerates the pace at which new drugs reach the market but also significantly reduces the resource and time costs involved in the research and development process[1]. For instance, doxycycline, a broad-spectrum antibiotic, was rapidly applied in clinical settings during the COVID-19 outbreak through drug repurposing, enhancing public health response efficiency and achieving notable results[2]. Thalidomide, once banned for its teratogenic effects, has been repurposed for treating multiple myeloma and other cancers, reducing drug development costs and improving patient survival rates[3]. Meanwhile, network pharmacology has provided a theoretical basis for identifying new drug indications by constructing and analyzing the interaction networks between drugs and their biological targets[4]. For example, the use of network pharmacology methods to analyze the interactions between known drugs and disease-related targets, thereby discovering the drugs’ potential new indications[5], or by analyzing the interaction networks between drug components to optimize drug combinations, enhancing therapeutic effects and reducing side effects. This method is particularly important for the treatment of complex diseases[6]. With the rapid development of bioinformatics and computational biology, there has been an increasing integration of advanced computational techniques such as deep graph models. These models are proving to be powerful tools in biomedical network analysis, thanks to their efficiency and accuracy in capturing complex network relationships, thus offering new insights into network pharmacology strategies[7].
@@
Deep graph models have brought revolutionary progress in the exploration of drug repurposing, particularly in the detailed revelation of chemical-gene interactions[8, 9]. Gaudelet et al. [10] explored the use of graph machine learning, including Graph Attention Networks (GATs), to predict drug-target interactions, offering new strategies for drug discovery. Zitnik et al. [11] used Graph Convolutional Networks (GCNs) to predict potential side effects of drug combinations, effectively enhancing the safety evaluation of polypharmacy by analyzing drug-drug interaction networks. Liang et al. [12] utilized GCNs to analyze the Drug-Gene Interaction database (DGIdb), showcasing the potential applications of GCNs in drug repurposing. However, despite their great performances, existing models still face several critical limitations. Models often fail to consider the sign of edges in the network[13, 14], such as whether the mechanism of actions by a drug is stimulatory or inhibitory.For example, in the DrugBank database, the relations between drugs and targets could be polar. Relations such as ’agonist’, ’activator’ and ’inducer’ indicate positive effects by the drug, whereas ’inhibitor’ , ’blocker’ and ’antagonist’ suggest negative effects. Often, terms like ’affect’ are also used, merely indicating an interaction. These nuances in biological significance are frequently ignored in unsigned networks, as models struggle to differentiate between labels that signify positive, negative, or neutral effects. The sign of edges (or polar edges/relations) is crucial for a deep understanding and accurate prediction of complex interactions within biological networks. The introduction of signed networks in protein-protein interaction network edge prediction by Mason et al.[15] and in gene co-expression networks by Kuhn et al.[16] has achieved significant success. Previous research has proven that constructing signed networks using graph convolutional neural networks for predicting interactions between biological entities can achieve excellent results. Teams led by Yang [17] and Chen[18] achieved notable outcomes in protein-protein interaction data and drug-drug interaction data, respectively. But the incorporation of mechanism of actions by drugs into the networks has not been studied yet. Moreover, the lack of uniformity in defining networks across various studies complicates the universality and comparability of research outcomes[19]. The field has yet to reach a consensus on determining the essential elements and information within networks, highlighting a significant gap in current research methodologies.
@@
To address these challenges, we have taken several innovative strategies. First, we developed a sophisticated deep graph model capable of handling signed networks. It not only addresses the limitation of existing methods that often ignore the polarity of the network, but is also compatible with non-polar relations[20]. In our model, the network contains polar relations (e.g. activation and inhibition) that are mutually exclusive to each other, and it also contains non-polar relations (e.g. binding) that could be co-existent. By identifying these complex and multifaceted relationships, our model provides a more detailed and dynamic perspective on chemical-gene interactions. We also implemented new evaluation metrics for the chemical-gene interaction network. Such metrics have been demonstrated to significantly improve the consistency and accuracy of model evaluations, as evident by their successful application in various advanced computational models[21]. At last, we conducted a thorough analysis of the characteristics of different components within the network and their impact on model performance. Based on this analysis, we proposed a series of strategies to optimize the model’s predictive capabilities. These strategies include data preprocessing, model structure optimization, and negative sampling during the training process, collectively aimed at enhancing the model’s accuracy and reliability.
@@ 2 Results
2.1 Experimental Setup
All experiments were conducted on a server equipped with an NVIDIA GeForce RTX 3090. The model’s encoder was a two-layer graph convolutional network optimized with Chebyshev polynomials[22], and the model was optimized using the Adam optimizer, with an early stopping mechanism added during the training process. The early stopping function is based on the model’s validation loss value, with a patience of 5. The final model was trained for 50 epochs with a batch size of 8192, and the best model parameters were found at epoch 48. The model was tested using a test size of 8192, and the test results were saved. The list of parameters used in the model is shown in the following Table 1.
Table 1: List of Model Training Parameters
Parameter Default Description
epoch 50 Number of epoches
hidden_dimensions [128, 64, 32, 16] List of hidden dimensions for layers
batch_size 8192 Batch size for training
test_batch True Whether to test the data per batch
test_batch_size 8192 Batch size for testing
print_step 10 Steps of printing information
average_precision_k 20 k for average precision at k
@@ 2.1 Experimental Setup
2.2 Performance Of Average Prediction
In this study, we first introduced a model that integrates the Relational Graph Convolutional Network with a Tensor Decomposition decoder, aimed at improving the accuracy of predicting interactions between genes and chemicals. We evaluated the model’s performance across multiple metrics, including the Area Under the Receiver Operating Characteristic curve (AUROC) and the Area Under the Precision-Recall Curve (AUPRC), as well as the average precision at the top 20 predictions (AP@20).We used three metrics to compare the prediction performances for gene-to-chemical ($P_{\mathrm{gene-chem}}$), chemical-to-gene ($P_{\mathrm{chem-gene}}$), and the average of both predictions ($P_{\mathrm{average}}$). The results (Table 2) indicate that $P_{\mathrm{average}}$ performs the best across all three metrics, with an AUROC of 0.966, AUPRC of 0.961, and AP@20 of 0.978. In contrast, $P_{\mathrm{chem-gene}}$ outperforms $P_{\mathrm{gene-chem}}$ in both AUROC and AUPRC, achieving scores of 0.939 and 0.925 respectively, compared to 0.882 and 0.841 for $P_{\mathrm{gene-chem}}$. In the AP@20 metric, $P_{\mathrm{chem-gene}}$ also shows superior predictive performance (0.824), significantly higher than that of $P_{\mathrm{gene-chem}}$ (0.689). These findings emphasize the superior overall performance of the average prediction model and also reveal potential differences in efficiency between different prediction directions.
Table 2: Model Training Results Metrics
Prediction $AUROC$ $AUPRC$ $AP@20$
$P_{\mathrm{gene-chem}}$ 0.882 0.841 0.689
$P_{\mathrm{chem-gene}}$ 0.939 0.925 0.824
$P_{\mathrm{average}}$ 0.966 0.961 0.978
@@ 2.2 Performance Of Average Prediction
The model’s ROC and PRC curves are shown in the following (Fig. 1). A and D shows results predicting genes to chemicals relations ($P_{\mathrm{gene-chem}}$) (Fig. 1 A,D); B and E shows the results predicting chemicals to genes relations ($P_{\mathrm{chem-gene}}$) (Fig. 1 B,E); C and F shows the average predictive performance of the $P_{\mathrm{average}}$ model on both tasks(Fig. 1 C,F).And these curves represent different polarity relationship edges respectively.
Figure 1: Model Training Results: A-C show the ROC curves for gene-to-chemical (A), chemical-to-gene (B), and average predictions (C). D-F display the corresponding PRC curves. Curve 1 represents to the ’Increase’ polarity relation edges. Curve 2 represents to the ’Decrease’ polarity relation edges. Curve 3 represents to the ’Binding’ and ’Affect’ relation edges.
The figures reveal that the $P_{\mathrm{average}}$ model rapidly ascends to nearly perfect true positive rates on the ROC curve while maintaining a low false positive rate, indicating its high effectiveness in distinguishing non-interactions between chemicals and genes. In the PRC curve, the $P_{\mathrm{average}}$ model also displays precision and recall rates that exceed those of these two single-link prediction models, especially maintaining high precision at high recall levels, indicating the $P_{\mathrm{average}}$ model’s significant advantage in ensuring the credibility of the prediction results. Overall, these graphs confirm the good performance of our model in prediction tasks.
2.3 Conflicting Edge Sampling Strategy
In our study, “edge conflicts” specifically refer to the conflicts during polarity predictions by the model, such as the same sample being predicted as both “activation” and “inhibition”. To avoid such conflicts, the model has to accurately distinguish between polar edges, such as activation and inhibition, which is crucial for understanding and predicting network behavior. By introducing the Cannot-Link (CL) conflicting edge sampling strategy, our model demonstrates enhanced discriminative capabilities when handling data with polar edges, significantly improving the accuracy in identifying polar relationships.
Although AUC and AP@k provide global view of model performances, they ignore the local intricacy when evaluating heterogenous networks. In our experiments, we explored the impact of the CL conflicting edge sampling strategy when classifying Must-Link (ML)-CL relationships using the RGCNTDDecoder model. Specifically, when applying the CL sampling strategy (YES), the AUROC value reached 0.959, the AUPRC value was 0.945, AP@20 was 0.902. In contrast, without using the CL sampling strategy (NO), the AUROC and AUPRC values were slightly higher, at 0.987 and 0.988 respectively, with AP@20 at 0.986. Further, considering the two new metrics we designed to evaluate the ability of handling polar relations, after using the CL sampling, the $AUC_{polarity}$ increased from 0.378 to 0.696 and the CP@500 also shown a considerable improvement from 0.766 to 0.862. Both metrics demonstrated that the model acquired the capability to distinguish polar relationships (Table 3). This indicates that although the CL sampling strategy may slightly reduce performance on some traditional metrics, it significantly lowers the probability of prediction conflicts when dealing with data of polar relations.
Table 3: Comparison of Model Performance Before and After Applying the CL Conflicting Edge Sampling Strategy
CL sampling $AUROC$ $AUPRC$ $AP@20$ $AUC_{polarity}$ $CP@500$
YES 0.985 0.985 1.000 0.696 0.862
NO 0.966 0.962 0.978 0.378 0.766

### v2

@@ Abstract
[…]
@@ 1 Introduction
In the new era of pharmaceutical research and development, the drug repurposing strategy plays a particularly critical role in exploring innovative therapies. This approach, through the rediscovery of existing drugs for new pathways or diseases, not only accelerates the pace at which new drugs reach the market but also significantly reduces the resource and time costs involved in the research and development process[1]. For instance, doxycycline, a broad-spectrum antibiotic, was rapidly applied in clinical settings during the COVID-19 outbreak through drug repurposing, enhancing public health response efficiency and achieving notable results[2]. Thalidomide, once banned for its teratogenic effects, has been repurposed for treating multiple myeloma and other cancers, reducing drug development costs and improving patient survival rates[3]. Meanwhile, network pharmacology has provided a theoretical basis for identifying new drug indications by constructing and analyzing the interaction networks between drugs and their biological targets[4]. For example, the use of network pharmacology methods to analyze the interactions between known drugs and disease-related targets, thereby discovering the drugs’ potential new indications[5], or by analyzing the interaction networks between drug components to optimize drug combinations, enhancing therapeutic effects and reducing side effects. This method is significant for the treatment of complex diseases[6]. With the rapid development of bioinformatics and computational biology, there has been an increasing integration of advanced computational techniques such as deep graph models. These models are proving to be powerful tools in biomedical network analysis, thanks to their efficiency and accuracy in capturing complex network relationships, thus offering new insights into network pharmacology strategies[7].
@@
Deep graph models have brought revolutionary progress in the exploration of drug repurposing, particularly in the detailed revelation of chemical-gene interactions[8, 9]. Gaudelet et al. [10] explored the use of graph machine learning, including Graph Attention Networks (GATs), to predict drug-target interactions, offering new strategies for drug discovery. Zitnik et al. [11] used Graph Convolutional Networks (GCNs) to predict potential side effects of drug combinations, effectively enhancing the safety evaluation of polypharmacy by analyzing drug-drug interaction networks. Liang et al. [12] utilized GCNs to analyze the Drug-Gene Interaction database (DGIdb), showcasing the potential applications of GCNs in drug repurposing. However, despite their great performances, existing models still face several critical limitations. Models often fail to consider the sign of edges in the network[13, 14], such as whether the mechanism of actions by a drug is stimulatory or inhibitory. For example, in the DrugBank database, the relations between drugs and targets could be polar. Relations such as ’agonist’, ‘activator’, and ’inducer’ indicate positive effects of the drug, whereas ’inhibitor’, ’blocker’, and ’antagonist’ suggest negative effects. Often, terms like ’affect’ are also used, merely indicating an interaction. These nuances in biological significance are frequently ignored in unsigned networks as models struggle to differentiate between labels that signify positive, negative, or neutral effects. The sign of edges (or polar edges/relations) is crucial for a deep understanding and accurate prediction of complex interactions within biological networks. The introduction of signed networks in protein-protein interaction network edge prediction by Mason et al.[15] and in gene co-expression networks by Kuhn et al.[16] has achieved significant success. Previous research has proven that constructing signed networks and using graph convolutional neural networks for predicting interactions between biological entities can achieve excellent results. Teams led by Yang [17] and Chen[18] achieved notable outcomes in protein-protein interaction data and drug-drug interaction data, respectively. However, incorporating the mechanism of actions by chemicals into the networks has not yet been studied.
Figure 1: To model a triplet pair within the network, the edges could be polar (e.g., positive or negative) or non-polar (e.g., neutral). In cases where polarity is present, the model is expected to accurately distinguish and identify the edge’s polarity.
@@
In modeling chemical-gene interaction networks, the current prediction models are designed solely for predicting interaction types without considering the polarity of these relationships. Additionally, they have not been trained to differentiate between polar and non-polar edges. We expect the model to exhibit a more significant probability difference for opposite polarities, as the model can discriminate between different types of edges by predicting distinguishing probabilities. In contrast, we do not anticipate models that predict very similar probabilities across various polarities, which are unable to distinguish between different types of edges (Fig.1). Although signed networks have been proposed as a solution to address polarity, these networks cannot accommodate the presence of non-polar edges (such as ‘binding’ or ‘affect’). This limitation highlights the need for models that can handle both polar and non-polar edges in a unified framework. To address these challenges, we have taken several innovative strategies. First, we developed a sophisticated deep graph model capable of handling signed networks. It not only addresses the limitation of existing methods that often ignore the polarity of the network but is also compatible with non-polar relations[19]. In our model, the network contains polar relations (e.g., activation and inhibition) that are mutually exclusive to each other, and it also contains non-polar relations (e.g., binding) that could be co-existent. By identifying these complex and multifaceted relationships, our model provides a more detailed and dynamic perspective on chemical-gene interactions. We also implemented new evaluation metrics for the chemical-gene interaction network. Such metrics have been demonstrated to significantly improve the consistency and accuracy of model evaluations[20]. At last, we conducted a thorough analysis of the characteristics of different components within the network and their impact on model performance. Based on this analysis, we proposed a series of strategies to optimize the model’s predictive capabilities. These strategies include data preprocessing, model structure optimization, and negative sampling during the training process, collectively aimed at enhancing the model’s accuracy and reliability.
@@ 2 Results
2.1 Performance Of Models Prediction
In this study, we designed and implemented a novel decoder model that integrates Relational Graph Convolutional Networks with tensor decomposition (RGCNTD), to enhance the accuracy of predicting chemical-gene interactions. We utilized several widely recognized classification metrics to comprehensively evaluate the model’s performance, including macro AUROC, micro AUROC, macro AUPRC, micro AUPRC, and the average precision of the top 20 predictions (AP@20). Leveraging the encoder framework provided by BioNet[21], we compared the performance of the original BioNet model with our RGCNTD model to validate its improvements. The results demonstrate that RGCNTD outperformed BioNet across all evaluation metrics, showcasing superior classification capability, enhanced robustness, and improved adaptability to imbalanced datasets.
Specifically, the RGCNTD model achieved scores of 0.966 and 0.980 in macro AUROC and micro AUROC, respectively, significantly surpassing the corresponding scores of 0.939 and 0.959 achieved by BioNet. This indicates a substantial improvement in both overall classification performance and the classification across all classes. In terms of macro AUPRC and micro AUPRC, the RGCNTD model attained scores of 0.961 and 0.971, respectively, further highlighting its superior handling of imbalanced data, particularly in identifying positive samples from minority classes. Moreover, the RGCNTD model excelled in the AP@20 metric, achieving a score of 0.968, a significant improvement over BioNet’s 0.824, demonstrating marked advancements in accuracy for the top 20 predictions. These results validate the RGCNTD model’s exceptional performance in predicting chemical-gene interactions.
To further substantiate the effectiveness of the RGCNTD model, we conducted a comparative analysis against several mainstream models(Table 1), including RGCN, GraphSAGE, and TransE. The results demonstrate that the RGCNTD model outperformed these alternatives across all metrics. The RGCN model achieved macro AUROC and micro AUROC scores of 0.530 and 0.572, respectively, reflecting significant limitations in overall classification performance and adaptability to class imbalance. GraphSAGE exhibited an even weaker performance, with a micro AUROC score of only 0.541. Although TransE showed marginally better performance than RGCN and GraphSAGE, its AUROC and AUPRC scores remained below 0.6, falling considerably short of the RGCNTD model. In the AP@20 metric, the RGCNTD model excelled with a score of 0.968, far surpassing the scores of 0.376 and 0.367 achieved by RGCN and GraphSAGE, respectively, which rendered their performance almost impractical for real-world applications.
Table 1: Model Performance Metrics
Prediction MACRO_AUROC MICRO_AUROC MACRO_AUPRC MICRO_AUPRC AP@20
RGCNTD 0.966 0.980 0.961 0.971 0.968
BioNet 0.939 0.959 0.925 0.937 0.824
RGCN 0.530 0.572 0.530 0.509 0.376
GraphSAGE 0.531 0.541 0.450 0.489 0.367
TransE 0.553 0.556 0.599 0.610 0.489
@@ 2.1 Performance Of Models Prediction
2.2 Conflicting Edge Sampling Strategy
In our study, ‘edge conflicts’ specifically refer to the conflicts during polarity predictions by the model, such as the same sample being predicted as both ‘activation’ and ‘inhibition’. To avoid such conflicts, the model has to accurately distinguish between polar edges, such as activation and inhibition, which is crucial for understanding and predicting network behavior.
We analyzed the conflicting predictions among RGCNTD, RGCN, GraphSAGE, and TransE models by examining the paired line plots shown in Fig.2 illustrate the performance of the four models in polar edge analysis. The left side of each figure represents ‘Decrease’ interactions, while the right side represents ‘Increase’ interactions. The lines connecting the log2-transformed rankings show how well the models distinguish between the two types of interactions. From these figures, we can observe that the RGCNTD (Fig.2 A) has the smallest overlap between the left and right sides, indicating its strong ability to differentiate edge polarity. However, there are still noticeable cases of polarity conflicts in its predictions. In contrast, the RGCN and the TransE models show more significant overlapping regions, highlighting the limitations of these models in distinguishing edge polarity, with conflicting predictions remaining widespread. The GraphSAGE model performs the worst in this regard, with almost complete overlap between the left and right sides, indicating its inability to distinguish between ‘Increase’ and ‘Decrease’ interactions effectively.
By introducing the Cannot-Link (CL) conflicting edge sampling strategy, our model demonstrates enhanced discriminative capabilities when handling data with polar edges, significantly improving the accuracy in identifying polar relationships. After introducing the conflict sampling strategy (CL sampling), we re-evaluated several models and observed performance improvements across various metrics(Table 2). Table 2 illustrates the performance differences of various models before and after introducing the CL sampling strategy. After applying CL sampling, the RGCNTD-CL showed significant improvements, with AUROC increasing from 0.966 to 0.985, AUPRC rising from 0.961 to 0.975, and AP@20 achieving a perfect score of 1. These results demonstrate that the CL sampling strategy greatly enhanced the generalization ability and accuracy of the RGCNTD model in recognizing complex interactions. For other baseline models, improvements were also observed with CL sampling, though not as pronounced as those for the RGCNTD model. RGCN-CL exhibited improvements in AP@20, which increased from 0.376 to 0.421, although AUROC declined from 0.530 to 0.509. GraphSAGE-CL showed more noticeable improvements, with macro AUROC increasing from 0.531 to 0.550 and AP@20 rising from 0.367 to 0.437, suggesting that the CL sampling strategy enhanced its performance. TransE-CL displayed slight improvements, with macro AUROC increasing from 0.553 to 0.573 and AP@20 rising from 0.489 to 0.510. Overall, the CL sampling strategy had the most significant impact on the RGCNTD model, while the improvements for other baseline models were also noticeable.
Table 2: Model Performance Metrics with and without CL
Model MACRO_AUROC MICRO_AUROC MACRO_AUPRC MICRO_AUPRC AP@20
RGCNTD 0.966 0.980 0.961 0.971 0.968
RGCNTD–CL 0.985 0.980 0.975 0.974 1
RGCN 0.530 0.572 0.530 0.509 0.376
RGCN–CL 0.509 0.547 0.533 0.512 0.421
GraphSAGE 0.531 0.541 0.450 0.489 0.367
GraphSAGE–CL 0.550 0.553 0.477 0.493 0.437
TransE 0.553 0.556 0.599 0.610 0.489
TransE–CL 0.573 0.569 0.599 0.611 0.510
@@ 2.2 Conflicting Edge Sampling Strategy
The impact of the CL sampling strategy on the model’s ability to distinguish polarity can be observed using the paired line plot in Fig.2. This plot visualizes the ranking of predictions for a pair of entities corresponding to ‘Increase’ and ‘Decrease’ polarity, allowing us to assess how effectively the model differentiates between them. For a given pair of entities, a well-performing model should assign distinctly different ranks to ‘Increase’ and ‘Decrease’. This means that in the paired line plot, a larger gap between the two ranks indicates better model performance. We compared our models with the baseline models. For the RGCNTD model (Fig.2 A), the overlapping region between the left and right sides significantly decreased, indicating that the CL sampling strategy effectively enhanced the model’s ability to differentiate between ”Increase” and ”Decrease” interactions, and significantly reduced edge conflicts. The RGCN model (Fig.2 B) and the TransE model (Fig.2 C) also showed improvements after the introduction of CL sampling, although conflicting regions remained relatively visible. The GraphSAGE model (Fig.2 D) exhibited the most significant improvement among all baseline models, with a substantial reduction in the overlap between the two sides, reflecting a marked enhancement in polarity distinction. Overall, the CL sampling strategy improved the ability to distinguish edge polarity across all models, with the RGCNTD model benefiting the most. This demonstrates the effectiveness of the strategy in addressing edge conflict issues.
Figure 2: The figures illustrate the performance of four models (RGCNTD, RGCN, GraphSAGE, and TransE) in polarity edge analysis. The left-side plots show the models’ performance without the CL sampling strategy, while the right-side plots display results after incorporating the strategy. The log2-transformed ranking lines connecting the two sides highlight each model’s ability to differentiate between the two types of interactions
2.3 Designed Metrics for Polairty Prediction
By evaluating two custom metrics (Table 3), AUC_polarity and CP@500, we could more accurately capture the models’ performance in polarity edge and high-confidence predictions. The results demonstrate that the conflict sampling strategy (CL sampling) brought significant improvements, especially for our model. After introducing CL sampling, the RGCNTD model’s AUC_polarity increased from 0.378 to 0.696, and CP@500 rose from 0.762 to 0.866. These improvements indicate that CL sampling greatly enhanced the model’s ability to differentiate edge polarity and more accurately predict interactions. These two metrics effectively validate the positive impact of CL sampling on handling complex interactions. For baseline models such as RGCN, GraphSAGE, and TransE, although CL sampling also improved AUC_polarity and CP@500, the effects were not as pronounced as with our model. For instance, RGCN’s AUC_polarity increased from 0.256 to 0.327, and CP@500 improved from 0.237 to 0.471, while the improvements for GraphSAGE and TransE were similarly limited. Despite some improvements in handling polarity conflicts with the help of CL sampling, the limited capacity of these baseline models restricted the extent of improvement.
This result suggests that our custom metrics, AUC_polarity and CP@500, effectively evaluate model performance in polarity prediction scenarios. Not only do they confirm the effectiveness of the CL sampling strategy, but they also highlight the significant advantage of the RGCNTD model in addressing complex interactions and polarity edge prediction.
Table 3: AUC Polarity and CP@500 Metrics
Model $AUC_{\text{polarity}}$ CP@500
RGCNTD 0.378 0.762
RGCNTD–CL 0.696 0.866
RGCN 0.256 0.237
RGCN–CL 0.327 0.471
GraphSAGE 0.212 0.218
GraphSAGE–CL 0.378 0.451
TransE 0.102 0.152
TransE–CL 0.410 0.479